Please help!
The sum of three numbers is 29. The greatest number is four times the smallest number. the remaining number is 1 more than twice the smallest number. Find the three numbers.

Answer by ankor@dixie-net.com(22740)

Three numbers: a, b, c; (a is the largest)
:
Write an equation for each statement
:
The sum of three numbers is 29.
a + b + c = 29
:
The greatest number is four times the smallest number.
a = 4c
:
the remaining number is 1 more than twice the smallest number.
b = 2c + 1
:
Find the three numbers.
using the 1st equation, replace a with 4c, replace b with (2c+1)
4c + (2c+1) + c = 29
7c + 1 = 29
7c = 29-1
c = 28/7
c = 4
then
a = 4(4) = 16
b = 2(4) + 1 = 9
:
16, 9, 4 are the three numbers
:
:
You can check the solutions by finding their sum